OpenStudy (anonymous):

The ratio of tagged fish to untagged fish was 2 to 9. Ninety fish were tagged. How many fish were untagged?

OpenStudy (texaschic101):

2/9 = 90/x 2x = 810 x = 810/2 x = 405 check.. 2/9 = 90/405 cross multiply (2)(405) / (9)(90) 810 = 810 (correct) if there is 90 tagged, then there is 405 untagged

OpenStudy (anonymous):

thank you!! @texaschic101

OpenStudy (texaschic101):

sure thing :)